OpenStudy (anonymous):

find all the real zeros of the polynomial function f(x)=49-x^2

OpenStudy (anonymous):

y = 49 - x^2 Because we are looking for "zeros", we are looking for all the points at which y=0. 0 = 49 - x^2 Let's multiply both sides by -1 for easier calculations, 0 = x^2 - 49 Add 49, 49 = x^2 x = sqrt[49] = +-7

OpenStudy (anonymous):

the answer is 7. :)